# Photodriven Mott insulating heterostructures: A steady-state study of impact ionization processes

### v1

@@ IV.3 Direct connection to wide-band metallic leads
where  [^14] Formally the WBL corresponds to taking $v_{\rho}\to\infty$, $\gamma_{\rho}\to\infty$ with finite $\Gamma_{\rho}$. $\Gamma_{\rho}\equiv 2v_{\rho}^{2}/\gamma_{\rho}$, while the Keldysh component in (21) is obtained by means of the fluctuation-dissipation theorem [43]. The WBL parameters are such that the magnitude of the gap and the bandwidth of the Hubbard bands are the same as in Sec. IV, i.e. $\Delta_{\text{pg}}\approx 3$ and $W_{\text{b}}\approx 8.5$.
Figure 10: (a) Photocurrent $j$ and (b) fluctuation of double occupancy $\Delta N_{\text{D},z}$ for the layer $z=2$, as function of the driving frequency $\Omega$ for selected values of the WBL rate $\Gamma$. (c) Occupation function $N_{z}(\omega)$ for the different layers at $\Omega=9$ for $\Gamma=0.16,0.02$. Dashed brown line in (a) represents the photocurrent $j$ for $\phi=0$ at $\Gamma=0.16$. Error bars corresponding to $\sigma_{j}$ in (a) are not shown for better visualization of the data. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
We start by analyzing the photocurrent $j$ as function of the driving frequency $\Omega$ for different values of the WBL rate $\Gamma=\Gamma_{\rho}$. In Fig. 10(a)  [^15] The curves relative to the $\Gamma=0.02$ for $\Omega<7$ exhibit some numerical instabilities due to the very small coupling to the leads. we observe that the peak of the photocurrent $j$ occurs at smaller driving frequencies $\Omega\approx 9$ and its magnitude is roughly one order of magnitude smaller than the results in Fig. 3(a)  [^16] Larger values of the driving frequency, i.e. $\Omega\gtrsim 14$, are responsible for a non-physical increase of $j$.. This is due to the mentioned effect that photoexcited carriers can escape to the ”wrong” side of the device. Moreover, the slight kink $\sim 2\Delta_{\textrm{pg}}$ in the $j$-$\Omega$ curve present in Fig. 3(a) is here completely absent at the corresponding $\Gamma=0.16$ (blue curve in Fig. 10(a)). For smaller values of the WBL rate $\Gamma$ there is also no evidence of such change of slope, which instead would be present with lower hybridizations in the setup discussed in Secs. IV.1 and IV.2, due to the presence of II. Considering also the large field amplitude $E_{0}=2$ chosen in this situation, this clearly points to a drastic reduction of II in this setup.
On the other hand, a kink around $\sim 2\Delta_{\textrm{pg}}$ can be observed in $\Delta N_{\text{D},z}$, see Fig. 10(b). This kink gets particularly pronounced for small $\Gamma$, and is connected with a higher occupation of the UHB of the heterostructure (Fig. 10(c)). This may favor II, although its signatures are absent in the photocurrent.

### v2

@@ IV.3 Direct connection to wide-band metallic leads
where  [^14] Formally the WBL corresponds to taking $v_{\rho}\to\infty$, $\gamma_{\rho}\to\infty$ with finite $\Gamma_{\rho}$. $\Gamma_{\rho}\equiv 2v_{\rho}^{2}/\gamma_{\rho}$, while the Keldysh component in (21) is obtained by means of the fluctuation-dissipation theorem [43]. The WBL parameters are such that the magnitude of the gap and the bandwidth of the Hubbard bands are the same as in Sec. IV, i.e. $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.
Figure 8: (a) Photocurrent $j$ and (b) fluctuation of double occupancy $\Delta N_{\text{D},z}$ for the layer $z=2$, as function of the driving frequency $\Omega$ for selected values of the WBL rate $\Gamma$. (c) Occupation function $N_{z}(\omega)$ for the different layers at $\Omega=9$ for $\Gamma=0.16,0.02$. Dashed brown line in (a) represents the photocurrent $j$ for $\phi=0$ at $\Gamma=0.16$. Error bars corresponding to $\sigma_{j}$ in (a) are not shown for better visualization of the data. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
We start by analyzing the photocurrent $j$ as function of the driving frequency $\Omega$ for different values of the WBL rate $\Gamma=\Gamma_{\rho}$. In Fig. 8(a)  [^15] The curves relative to the $\Gamma=0.02$ for $\Omega<7$ exhibit some numerical instabilities due to the very small coupling to the leads. we observe that the peak of the photocurrent $j$ occurs at smaller driving frequencies $\Omega\approx 9$ and its magnitude is roughly one order of magnitude smaller than the results in Fig. 3(a)  [^16] Larger values of the driving frequency, i.e. $\Omega\gtrsim 14$, are responsible for a non-physical increase of $j$.. This is due to the mentioned effect that photoexcited carriers can escape to the ”wrong” side of the device. Moreover, the slight kink $\sim 2\Delta_{\textrm{g}}$ in the $j$-$\Omega$ curve present in Fig. 3(a) is here completely absent at the corresponding $\Gamma=0.16$ (blue curve in Fig. 8(a)). For smaller values of the WBL rate $\Gamma$ there is also no evidence of such change of slope, which instead would be present with lower hybridizations in the setup discussed in Secs. IV.1 and IV.2, due to the presence of II. Considering also the large field amplitude $E_{0}=2$ chosen in this situation, this clearly points to a drastic reduction of II in this setup.
On the other hand, a kink around $\sim 2\Delta_{\textrm{g}}$ can be observed in $\Delta N_{\text{D},z}$, see Fig. 8(b). This kink gets particularly pronounced for small $\Gamma$, and is connected with a higher occupation of the UHB of the heterostructure (Fig. 8(c)). This may favor II, although its signatures are absent in the photocurrent.